The velocity of a 2kg object increases from 2.0m/s to 4.0m/s by a constant force. Find the change in momentum of the object. Show complete solution.


1
Expert's answer
2022-01-10T16:55:51-0500

We can find the change in momentum of the object as follows:


"\Delta p=m\Delta v=m(v_f-v_i),""\Delta p=2\ kg\times(4\ \dfrac{m}{s}-2\ \dfrac{m}{s})=4\ \dfrac{kg\times m}{s}."